# Computational design of NDR tunnel diodes with ultra-high peak-to-valley current ratio based on two-dimensional cold metals: The case of NbSi2N4/HfSi2N4/NbSi2N4 lateral heterojunction diode

### v2

@@ I Introduction
Existing semiconductor-based NDR tunnel diodes suffer from low peak-to-valley current ratio (PVCR) values, which is attributed to the band tails tunneling that originates from the strong doping and dopant fluctuations [23, 24, 25]. To address this issue, we recently proposed a new semiconductor-free NDR tunnel diode concept with ultra-high PVCR [26]. Our proposed diode consists of two cold metal electrodes separated by a thin insulating tunnel barrier (see Fig. 1). The NDR effect arises from the unique electronic band structure of the cold-metal electrodes (see Fig. 2). A cold metal is obtained when the metallic, i.e. partially filled, bands are separated from all energetically higher and lower-lying bands by sufficiently large energy gaps. Specifically, the width of the isolated metallic bands around the Fermi level, as well as the energy gaps separating higher and lower-lying bands, determine the current-voltage (-) characteristics and PVCR of the tunnel diode. By choosing the appropriate cold metal electrode materials, either a conventional N-type or -type NDR effect can be achieved as sketched in Fig. 1(b,c).
@@ II Computational Method
The computational design of the presented lateral heterojunction tunnel diodes is based on density functional theory (DFT) within the QuantumATK T-2022.03 package [32]. We used a linear combination of atomic orbitals (LCAO) basis set (double zeta polarized) with norm-conserving FHI pseudopotentials [33]. The exchange-correlation functional was represented by a generalized gradient approximation (GGA) with Perdew-Burke-Ernzerhof (PBE) flavor[34]. The self-consistent DFT calculations are performed using a k-point grid with a density mesh cut-off of and the total energy converged to at least . To prevent interactions between the periodically repeated images, of vacuum were added and Dirichlet and Neumann boundary conditions are employed. The transport calculations were performed using DFT combined with a nonequilibrium Green’s function method (NEGF). We use a () -point grid in self-consistent DFT-NEGF calculations of the lateral tunnel diodes along the armchair (zigzag) orientation. The - characteristics were calculated within a Landauer approach [35], where . Here denotes the bias voltage, are the transmission coefficients and and are the Fermi-Dirac distributions of the left and right electrodes, respectively. The transmission coefficients for the lateral tunnel diodes along the armchair (zigzag) orientation are calculated using a ( ) -point grid and convergence tests can be found in the supplemental material [36].
@@ III Results and Discussion
In Fig. 5, we present the calculated - curves for tunnel diodes with tunnel barrier widths of and , for both armchair (black lines) and zigzag (red lines) directions. A maximum bias voltage of was chosen. We report a conventional -type NDR effect in both transport directions with a valley voltage of approximately , which is primarily set by the metallic band width of the cold metal electrodes (see Table 1). Besides this similarity, the - curves of the armchair and zigzag directions exhibit distinct behavior. In the armchair case, the peak current (cf. Fig. 2 )is achieved at very low bias voltages of about , while in the zigzag case it appears at higher bias voltages of >. These details can be directly attributed to the overlap of contributing electronic states of the left and right cold metal electrodes in different areas of the Brillouin zone. The mentioned overlap differs for armchair and zigzag-oriented tunnel diodes resulting in the presented - curves in Fig. 5. In a first approximation, the spin-independent transmission through an insulating barrier of width at zero bias is [39]. Here and are the interface transmission functions of the left and right metallic leads, respectively. Assuming perfect symmetry matching of the wave functions at the interfaces, the product coincides with the joint density of states of left and right electrodes. The remaining factor involves the exponential decay length of the wave function within the insulating barrier. The latter is directly connected to the complex band structure of the barrier, which will be discussed later in the letter. We introduce further insights on the specific structure of the - curves maxima via the energy- and state-resolved transmission, as well as projected interface band structures in the supplementary material [36].
@@ Acknowledgements.
- XYZ [2023] XYZ, Supplemental material, Phys. Rev. Applied 1, 000000 (2023).

### v3

# Computational design of tunnel diodes with negative differential resistance and ultrahigh peak-to-valley current ratio based on two-dimensional cold metals: The case of NbSi2N4/HfSi2N4/NbSi2N4 lateral heterojunction diode

@@ I Introduction
Existing semiconductor-based NDR tunnel diodes suffer from low peak-to-valley current ratio (PVCR) values, which is attributed to the band tails tunneling that originates from the strong doping and dopant fluctuations [23, 24, 25]. To address this issue, we recently proposed a new semiconductor-free NDR tunnel diode concept with ultra-high PVCR [26]. Our proposed diode consists of two cold metal electrodes separated by a thin insulating tunnel barrier (see Fig. 1). The NDR effect arises from the unique electronic band structure of the cold-metal electrodes (see Fig. 2). A cold metal is obtained when the metallic, i.e. partially filled, bands are separated from all energetically higher and lower-lying bands by sufficiently large energy gaps. Specifically, the width of the isolated metallic bands around the Fermi level, as well as the energy gaps separating higher and lower-lying bands, determine the current-voltage (-) characteristics and PVCR of the tunnel diode. By choosing the appropriate cold metal electrode materials, either a conventional -type or -type NDR effect can be achieved as sketched in Fig. 1(b,c).
@@ II Computational Method
The computational design of the presented lateral heterojunction tunnel diodes is based on density functional theory (DFT) within the QuantumATK T-2022.03 package [32]. We used a linear combination of atomic orbitals (LCAO) basis set (double zeta polarized) with norm-conserving FHI pseudopotentials [33]. The exchange-correlation functional was represented by a generalized gradient approximation (GGA) with Perdew-Burke-Ernzerhof (PBE) flavor[34]. The self-consistent DFT calculations are performed using a k-point grid with a density mesh cut-off of and the total energy converged to at least . To prevent interactions between the periodically repeated images, of vacuum were added and Dirichlet and Neumann boundary conditions are employed. The transport calculations were performed using DFT combined with a nonequilibrium Green’s function method (NEGF). We use a () -point grid in self-consistent DFT-NEGF calculations of the lateral tunnel diodes along the armchair (zigzag) orientation. The - characteristics were calculated within a Landauer approach [35], where . Here denotes the bias voltage, are the transmission coefficients and and are the Fermi-Dirac distributions of the left and right electrodes, respectively. The temperature was kept at throughout all calculations. The transmission coefficients for the lateral tunnel diodes along the armchair (zigzag) orientation are calculated using a ( ) -point grid and convergence tests can be found in the supplemental material [36].
@@ III Results and Discussion
In Fig. 5, we present the calculated - curves for tunnel diodes with tunnel barrier widths of and , for both armchair (black lines) and zigzag (red lines) directions. A maximum bias voltage of was chosen. We report a conventional -type NDR effect in both transport directions with a valley voltage of approximately , which is primarily set by the metallic band width of the cold metal electrodes (see Table 1). Besides this similarity, the - curves of the armchair and zigzag directions exhibit distinct behavior. In the armchair case, the peak current (cf. Fig. 2 )is achieved at very low bias voltages of about , while in the zigzag case it appears at higher bias voltages of . These details can be directly attributed to the overlap of contributing electronic states of the left and right cold metal electrodes in different areas of the Brillouin zone. The mentioned overlap differs for armchair and zigzag-oriented tunnel diodes resulting in the presented - curves in Fig. 5. In a first approximation, the spin-independent transmission through an insulating barrier of width at zero bias is [39]. Here and are the interface transmission functions of the left and right metallic leads, respectively. Assuming perfect symmetry matching of the wave functions at the interfaces, the product coincides with the joint density of states of left and right electrodes. The remaining factor involves the exponential decay length of the wave function within the insulating barrier. The latter is directly connected to the complex band structure of the barrier, which will be discussed later in the letter. We introduce further insights on the specific structure of the - curves maxima via the energy- and state-resolved transmission, as well as projected interface band structures in the supplementary material [36].
@@ Acknowledgements.
- [36] See Supplemental Material at [URL-will-be-inserted-by-publisher] for further insights into the characteristics of the devices.
@@
Supplemental Material: Computational design of tunnel diodes with negative differential resistance and ultrahigh peak-to-valley current ratio based on two-dimensional cold metals: The case of NbSi2N4/HfSi2N4/NbSi2N4 lateral heterojunction diode
P. Bodewei1, E. Şaşıoğlu1, N. F. Hinsche1, I. Mertig1 1Institute of Physics, Martin Luther University Halle-Wittenberg, 06120 Halle (Saale), Germany
(a) barrier
(b) barrier
In Fig. 13 the --curves for positive and negative bias are shown. While the --curve for the zigzag device with and barrier thickness are perfectly antisymmetric, we see that the --curves for devices in armchair direction show a slight asymmetry. The reason for that is the interface potential, building up between NbSi2N4 and HfSi2N4. Geometrically, we can see that the termination of the interface between the right electrode and left electrode with the barrier is different in armchair devices. Thus, a different interface potential at both interfaces builds up, leading to different current densities for positive and negative bias voltages, respectively.
(a) armchair device with tunnel barrier
(b) zigzag device with tunnel barrier
In Fig. 16 we see the transmission at different -points for different energies. Firstly, we see that there is no certain -point dominating the transmission, i.e. the transmission is distributed throughout the whole range of -points for both the armchair and the zigzag device. Further, we see that for higher energies of about we see a strong contribution of the transmission at the -point.
Bias
Bias
Bias
Bias
Bias
Bias
In Fig. 20 and Fig. 24 the - and energy-resolved transmission spectrum is shown for three different bias voltages. Each black curve represents the transmission for a single -point, and the red lines show the bias window (). In the following, we will call the transmission and the -resolved transmission. Since we are calculating the current within the Landauer-Büttiker approach, the main quantity is the transmission, which can be calculated by integrating over all -points. This is shown by the blue curve. Finally, one obtains the --characteristics by integrating over a broad energy range. Here, we show - and -resolved transmissions for the -barrier armchair (Fig. 20) and the -barrier zigzag device (Fig. 24). The bias voltages of , , and were chosen, as the maximum of the current in the armchair device occurs at and the maximum of the zigzag around . For the armchair device, we deduce that the area under the transmission curve is significantly higher at than for , underpinning the maxima in the --curves (fig. 13). For the zigzag device, we spot an increase in the area under the transmission curve for the higher bias voltages. Furthermore, one notices that the areas under the transmission curves for the armchair device are much smaller than for the zigzag device, supporting the difference in the current densities absolute values for both devices.